the sum of three consecutive integers is 1872
Answer by oberobic(2304)

Define 3 consecutive integers as:
x
x+1
x+2
.
x +x+1 +x+2 = 1872
3x + 3 = 1872
3x = 1869
x = 623
x+1 = 624
x+2 = 625
.
Answer: The three consecutive integers are 623, 624, and 625.
.
Done.